OpenStudy (anonymous):

Use the discriminant to determine all values of k that would result in the equation x^2 -kx+4=0 having equal roots

OpenStudy (watchmath):

\(D=k^2-4(4)=0\) \(k^2=16\) \(k=\pm 4\)